# Filtered Direct Preference Optimization

### v2

@@ B.2 Additional results
Figure 7 provides the learning curves for DPO and fDPO with the the 160M-sized LM, corresponding to the final performances depicted in Figure 1 (B) of the main text. The curves show that even though fDPO processes double the number of epochs compared to DPO, the total number of steps for fDPO is fewer than that for DPO. This reduction is due to the filtering process of fDPO, which decreased the data over epochs, resulting in fewer steps per epoch, as demonstrated in Figure 6. Additionally, when assessed using KL divergence, the performance of fDPO shows a trend towards converging with the DPO trained on the high-quality dataset, suggesting that fDPO can close the performance gap even when trained on mixed-quality data.
@@
Figure 7: The learning curves for DPO and fDPO using the 160M-sized LM on the mix-quality dataset. The horizontal axes of the two figures represent the number of training steps and the KL divergence with the initial LM, respectively. Normalized gold rewards are adjusted from 0 (initial) to 1 (DPO on high-quality).”
@@
Figure 8: The learning curves for DPO and fDPO using the 1.4B LM on the mix-quality dataset (left) and the 160M LM on the low-quality dataset (right), respectively. Normalized gold rewards are adjusted from 0 (initial) to 1 (DPO).”

### v3

@@ C.3.1 Detailed evaluation
Mix 4 0.496 -1.231 -0.367
Mix 5 0.495 -1.234 -0.370
Table 4: The evaluation scores of gold reward for AlpacaFarm dataset
Figure 6 provides the learning curves for DPO and fDPO with the the 160M-sized LM, corresponding to the final performances depicted in Figure 1 (B) of the main text. The curves show that even though fDPO processes double the number of epochs compared to DPO, the total number of steps for fDPO is fewer than that for DPO. This reduction is due to the filtering process of fDPO, which decreased the data over epochs, resulting in fewer steps per epoch, as demonstrated in Figure 8. Additionally, when assessed using KL divergence, the performance of fDPO shows a trend towards converging with the DPO trained on the high-quality dataset, suggesting that fDPO can reduce the performance gap even when trained on mixed-quality data.
@@
Figure 6: The learning curves for DPO and fDPO using the 160M-sized LM on the mix-quality dataset of AlpacaFarm. The horizontal axes of the figures represent the number of training steps and the KL divergence with the initial LM (SFT model), respectively, where the gold rewards are adjusted so that the average reward of the SFT model is zero.
@@
Figure 7: The learning curves for DPO and fDPO using the 1.4B LM on the mix-quality AlpacaFarm dataset (left) and the 160M LM on the low-quality AlpacaFarm dataset (right), respectively, where the gold rewards are adjusted so that the average reward of the SFT model is zero.
C.3.2 Analysis of filtered samples of fDPO
We examined how data was selectively discarded throughout the learning process of fDPO with the mix-quality dataset. Figure 8 presents the unfiltered ratio, accuracy, precision, and recall at each epoch. The unfiltered ratio reflects the proportion of data that remains after filtering. Accuracy reflects the overall correctness of the filtering decisions, both for deletion and retention of samples, based on their gold reward quality. Precision measures how accurately the samples decided for deletion were actually of lower quality, while recall evaluates the success in identifying and discarding all samples that warranted removal. The result of the unfiltered ratio indicates an exponential decay in the number of samples used in each epoch. The consistency of accuracy and precision across epochs suggests that data was discarded with a constant efficiency. The lower precision compared to accuracy can be attributed to the relatively small number of samples that warranted removal. Conversely, recall decreases with progressing epochs. This decline can be tied to the static errors within the proxy RM, leading to consistently overestimated $y_{c}$ samples, thus increasing their relative proportion over time. The figure contrasts various margin settings with the no-margin condition ($\epsilon=0$), revealing that larger margins lead to slower filtering speeds. Notably, as the margin increases, precision improves at the expense of recall. This trade-off indicates the importance of carefully tuning the margin parameter $\epsilon$ to balance filtering efficacy.
Figure 8: Unfiltered ratio, accuracy, precision, and recall throughout epochs in fDPO, comparing the effects of no margin ($\epsilon=0$) with various margin levels ($\epsilon>0$) on the filtering condition. Larger margins lead to slower filtering speeds but improve precision at the expense of recall, highlighting the need to carefully tune the margin parameter $\epsilon$.
C.4 Additional results on Anthropic HH Datasets
C.4.1 Detailed evaluation
The evaluation scores of the gold reward model for our preference datasets (original, SFT-model-generated, mix-quality) of the Anthropic HH dataset are detailed in Table 5. The mix-quality datasets (Mix 1-3) each consist of 25% randomly sampled responses from the original Anthropic HH dataset and 75% from the responses generated by the SFT model, using random seeds 1-3, respectively.
Dataset Type Chosen Mean Rejected Mean Overall Mean
Helpful Original -0.294 -1.549 -0.922
SFT Generated -0.613 -1.931 -1.272
Mix 1 -0.537 -1.836 -1.187
Mix 2 -0.532 -1.839 -1.185
Mix 3 -0.536 -1.833 -1.185
Harmless Original -3.142 -4.622 -3.882
SFT Generated -4.164 -5.455 -4.810
Mix 1 -3.905 -5.245 -4.575
Mix 2 -3.907 -5.250 -4.579
Mix 3 -3.914 -5.250 -4.582
Table 5: The evaluation scores of gold reward for Anthropic HH datasets.
Figure 9 provides using the helpful dataset and the harmless dataset of the Anthropic HH datasets, respectively. The curves show that even though fDPO processes double the number of epochs compared to DPO, the total number of steps for fDPO is fewer due to the filtering process.
(a) Helpful
(b) Harmless
Figure 9: The learning curves for DPO and fDPO using helpful dataset (left) and harmless dataset (right) of Anthropic HH datasets, respectively, where the gold rewards are adjusted so that the average reward of the SFT model is zero.
C.4.2 Future Evaluation
In addition to the aforementioned experiment, we conduct an experiment with 50% original and 50% SFT-generated responses. The detailed results of this experiment are provided in Table 6.
Dataset Method Gold RM Score (SFT=$0.0$) $\uparrow$ GPT-4o Evaluation (win rate vs. SFT) $\uparrow$
Helpful DPO 1.72 $\pm$ 0.03 0.575 $\pm$ 0.007
fDPO 1.85 $\pm$ 0.05 0.602 $\pm$ 0.010
Harmless DPO 2.74 $\pm$ 0.07 0.856 $\pm$ 0.012
fDPO 3.23 $\pm$ 0.01 0.955 $\pm$ 0.007
Table 6: Evaluation on the Anthropic HH datasets, where the responses consist of 50% original and 50% SFT-generated responses. The values represent the mean and standard error over 3 seeds.